# Dome-DETR: DETR with Density-Oriented Feature-Query Manipulation for Efficient Tiny Object Detection

### v1

@@ Abstract.
[…]
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: –; –; –isbn: 978-1-4503-XXXX-X/2018/06submissionid: 846ccs: Computing methodologies Object detection
Figure 1. Comparisons with other detectors in accuracy and GFLOPs on the AI-TOD-V2 test set. Circle size indicates model size. More detailed analysis can be found in Table 1.
@@ 1. Introduction
Object detection is a fundamental task in computer vision, underpinning a wide range of real-world applications such as autonomous driving and robotic navigation. In recent years, CNNs have driven significant advancements in object detection (Dai et al., 2016; Redmon et al., 2016; He et al., 2017; Lin et al., 2017; Liu et al., 2016; Girshick, 2015; Ren et al., 2016; Wang et al., 2023; Tang et al., 2024; Peng et al., 2024b). Meanwhile, the introduction of Detection Transformers (DETR) revolutionized the field by leveraging Transformer architectures for end-to-end detection (Carion et al., 2020). Subsequent variants, including Deformable-DETR (Zhu et al., 2020), DINO-DETR (Zhang et al., 2022) and D-FINE (Peng et al., 2024a), etc. (Chen et al., 2024; Liu et al., 2022; Meng et al., 2021; Huang et al., 2024b; Zhao et al., 2024) have further enhanced performance and speed. However, these advancements primarily target generic object detection, leaving critical challenges unresolved for detecting tiny objects—particularly in aerial imagery from drones or satellites.
@@
Tiny object detection, which aims to localize and classify objects occupying only a few pixels, remains a particularly difficult problem in computer vision. It is crucial for applications such as remote sensing, drone-based monitoring, and autonomous navigation. The primary challenge stems from the fragile and sparse feature representation of tiny objects. These objects heavily rely on low-level spatial details, which are often lost in deeper layers of feature hierarchies. Maintaining high-resolution feature maps help preserve such information but comes at the cost of increased computational complexity and memory consumption. Deformable DETR (Zhu et al., 2020) alleviates some of this burden through sparse deformable attention, yet still suffers from high inference latency due to its broad multi-scale attention. RT-DETR (Zhao et al., 2024) improves inference speed by decoupling intra-scale and cross-scale interactions but relies heavily on deep, low-resolution features, resulting in degraded performance on small-scale targets. These issues highlight the pressing need for detection frameworks that can balance fine-grained feature preservation with computational efficiency.
@@
Beyond effective feature representation, query allocation introduces another bottleneck for tiny object detection. Aerial images often exhibit high instance density with complex object distributions (Wang et al., 2021b; Zhu et al., 2021). This variability complicates the allocation of queries in object detection frameworks. Existing DETR-like methods (Carion et al., 2020; Zhu et al., 2020; Liu et al., 2022; Li et al., 2022; Zhang et al., 2022; Zhao et al., 2024) use a fixed number of queries ($K=100$ in DETR, $K=300$ in Deformable-DETR), limiting adaptability to object count variations. This simplifies implementation but reduces recall in dense scenes and wastes resources in sparse ones. For instance, aerial datasets such as AI-TOD-V2 (Wang et al., 2021b) can contain images with over 1,500 tiny objects—far surpassing the query capacity of conventional DETR variants. To address this, DDQ-DETR (Zhang et al., 2023) increases query density (using $K=900$) and applies class-agnostic NMS with manually set IoU thresholds to filter redundant predictions. However, its rigid query count and fixed NMS thresholds introduce low recall in dense scenes and remain insensitive to variations in instance density. Meanwhile, DQ-DETR (Huang et al., 2024a) introduces dynamic query adjustment through a categorical counting module that estimates a classified number for query allocation. Although promising, its counting head relies on manually tuned classification hyperparameters across different datasets, and integrating density prediction with feature enhancement modules leads to substantial computational overhead. These challenges underscore the need for an adaptive query mechanism that dynamically aligns query density with instance distributions while eliminating manual tuning and preserving efficiency. To this end, tiny object detection remains challenging due to the fragile nature of small object features and the inefficiency of fixed-query mechanisms in handling diverse instance densities.
In this paper, we introduce Dome-DETR, a novel end-to-end object detection framework which improves small and tiny object detection with minimal additional computation. First, in satellite and drone imagery—as well as in natural scenes—the foreground occupies only a small portion of the frame, as is shown in Figure 2. This results in excessive computation on background regions, which contain far less useful information for object detection. The observation above reveals the potential to accelerate the detection by focusing more on the foreground area. Besides, it is very intuitive that the shallow feature maps extracted from CNN backbones are rich in spatial density information. To exploit this motivation, we propose a lightweight (A) Density-Focal Extractor (DeFE), which generates density response heatmaps closely aligned with ground-truth annotations. These heatmaps can simultaneously enhance multi-scale features in the encoder and facilitate progressive query initialization in the decoder. Second, while deep features capture scene-level semantics, the instance-level morphological details present in shallow features are critical for accurate detection of small objects. However, their high resolution incurs substantial computational costs in attention mechanisms. To address this issue, we leverage density response maps from (A) to mask non-critical regions, enabling (B) Masked Window Attention Sparsification (MWAS) to focus only on significant windows. Lastly, conventional DETR’s fixed query mechanism struggles with the extreme object count variations typical of aerial images. Therefore, we introduce (C) Progressive Adaptive Query Initialization (PAQI), which allocates queries adaptively by decoding density responses into thresholds of dynamic bounding box suppression. This eliminates hand-crafted hyperparameters in previous methods while improving recall in dense scenes. Furthermore, we streamline computationally intensive components in existing real-time DETR architectures, striking a better balance between speed and accuracy.
@@
Figure 2. (A) Comparison of foreground proportions across remote sensing, drone imagery, and COCO datasets; and (B) Visualization of foreground regions (highlighted in yellow) in samples from these three datasets.
@@
Experimental results on the AI-TOD-V2 (Wang et al., 2021b) dataset demonstrate that our Dome-DETR achieves state-of-the-art performance in tiny object detection, surpassing existing models in both accuracy and efficiency. Specifically, our Dome-DETR-M and Dome-DETR-L models achieve 34.0% (+3.2 AP) and 34.6% (+3.3 AP), respectively, on the AI-TOD-V2 test set, while incurring computational costs of only 252.6 and 358.7 GFLOPs. We also achieve excellent results on the VisDrone (Zhu et al., 2021) val set, with Dome-DETR-L attaining 39.0% (+2.5 AP). By effectively addressing key challenges in feature representation and query manipulation, Dome-DETR represents a significant advancement in tiny object detection using DETR-based architectures, paving the way for future research in this area. In summary, our main contributions are as follows:
@@
  • We achieve state-of-the-art performance on the AI-TOD-V2 and VisDrone-DET-2019 datasets. Specifically, Dome-DETR-L achieves 34.6% AP on the AI-TOD-V2 test set and 39.0% AP on the VisDrone val set while maintaining a low computational cost, outperforming all existing state-of-the-art models.
@@ 3.1. Overview
As shown in Figure 3, our study proposes Dome-DETR, which is built upon the architecture of D-FINE (Peng et al., 2024a). We enhance the model with three components, i.e. (A) Density-Focal Extractor (DeFE) for density prediction, (B) Masked Window Attention Sparsification (MWAS) for efficient shallow feature enhancement, and (C) Progressive Adaptive Query Initialization (PAQI) for dynamic query manipulation.
@@ 3.2. Density-Focal Extractor
In DETRs, backbone generate spatially-channel encoded feature maps through multi-scale feature extraction. This mapping is established via end-to-end training, where the network learns to transform elementary visual features into high-level semantic representations. Consequently, foreground regions containing objects exhibit distinctive activation patterns, while background regions maintain low-response characteristics.
Besides, previous studies have proved that shallow features preserve essential spatial details for tiny objects (Ning et al., 2023; Tu et al., 2022), but they are hard to utilize in two reasons: (1) High-resolution feature maps induce prohibitive computational costs, and (2) Redundant background regions dominate spatial attention, diluting critical instance-level information.
To address this, we propose the Density-Focal Extractor (DeFE)—a lightweight module that explicitly learns instance density distributions to guide efficient feature enhancement and query allocation.
As illustrated in Figure 3(A), DeFE processes the shallowest backbone feature map $F_{S}\in\mathbbm{R}^{H\times W\times C}$through an optimized cascaded network, which employs depthwise separable convolutions with varying dilation rates (1, 2, 3) to capture multi-scale contextual information while maintaining efficiency. A lightweight attention mechanism then highlights salient regions by channel-wise feature recalibration. The processed features pass through a density prediction head comprising a 3×3 convolution and bilinear upsampling, generating a normalized density heatmap $D_{pred}\in\mathbbm{R}^{H\times W\times 1}$. Together, DeFE can be formulated as:
@@
where $f_{\mathrm{DSConv}}(\cdot)$ applies cascaded depthwise separable convolutions with different dilation rates to extract spatially rich features. $F_{G}$ is a compact global representation pooled from the feature map using Global Average Pooling(GAP). $W_{P}$ means a convolution layer that projects features into a single-channel density map, followed by sigmoid function and bilinear upsampling to match the original resolution and data distribution.
The ground-truth density map $D_{gt}$ is generated by convolving Gaussian kernels centered at each object’s center coordinates, with kernel sizes proportional to bounding box dimensions. This design encodes both instance locations and relative scales into a continuous supervision signal. To train DeFE, we propose the Density Recall Focal Loss (DRFL) that prioritizes accurate density estimation in critical regions:
(4) $\mathcal{L}_{DRFL}=\sum_{i,j}^{H,W}\left[\alpha_{i,j}(d_{i,j}^{pred}-d_{i,j}^{gt})^{2}+\beta\cdot\mathbbm{1}(d_{i,j}^{pred}<d_{i,j}^{gt})\cdot d_{i,j}^{gt}\right],$
where $\alpha_{i,j}=\sqrt{d_{i,j}^{gt}}$ adaptively weights positions based on ground-truth density intensity, and $\beta$ penalizes underestimation in high-density regions. This formulation ensures balanced learning across sparse and dense areas while preventing FN in crowded scenes. While maintaining excellent density estimation accuracy, the entire module adds only 0.8M parameters due to optimized depthwise convolutions and attention mechanisms.
@@
3.3. Masked Window Attention Sparsification
The high resolution of shallow features is essential for capturing fine-grained details of tiny objects, but it also introduces substantial computational burden. Processing these dense feature maps with global attention mechanisms results in excessive memory consumption and latency. To mitigate this, we propose Masked Window Attention Sparsification (MWAS), which selectively focuses computational resources on critical foreground regions while discarding redundant background information. As illustrated in Figure 3(B), MWAS operates in two key stages: Foreground Token Pruning and Axis Permuted Encoder (APE).
Firstly, by using the predicted density map from Density-Focal Extractor (DeFE), we generate a binary mask where high-density regions are retained, and low-density background regions are pruned. This mechanism significantly reduces the number of tokens that participate in attention computation while preserving essential object details.
Density-Based Mask Generation. Given the shallow feature map $F_{S}$ extracted from the backbone and density heatmap $D_{\mathrm{pred}}\in\mathbbm{R}^{H\times W\times 1}$ from DeFE, where each element $d_{i,j}^{\mathrm{pred}}$ represents the estimated object density at spatial position $(i,j)$. The binary mask $M_{b}\in\{0,1\}^{H\times W}$ is generated by thresholding $D_{\mathrm{pred}}$ with an adaptive threshold $T_{b}$:
@@ 3.3. Masked Window Attention Sparsification
5:  $\mathbf{A}\leftarrow\text{GenerateAnchors}(\mathbf{S})$
6:  $\mathbf{M}_{out},\mathbf{S}_{enc}\leftarrow\text{ClsHead}(\mathbf{M})$
@@
8:  $\mathbf{M}_{T},\mathbf{S}_{T},\mathbf{A}_{T}\leftarrow\text{SelectTopK}(\mathbf{M}_{out},\mathbf{S}_{enc},\mathbf{A},K_{M})$
@@
10:  $\mathbf{M}_{1},\mathbf{S}_{1},\mathbf{A}_{1}\leftarrow\mathbf{M}_{T}[:K_{N}],\mathbf{S}_{T}[:K_{N}],\mathbf{A}_{T}[:K_{N}]$
11:  $\mathbf{M}_{2},\mathbf{S}_{2},\mathbf{A}_{2}\leftarrow\mathbf{M}_{T}[K_{N}:K_{M}],\mathbf{S}_{T}[K_{N}:K_{M}],\mathbf{A}_{T}[K_{N}:K_{M}]$
@@
13:  $\mathbf{A}_{2}\leftarrow\mathbf{A}_{2}[\mathbf{W}[\mathbf{A}_{2}]>0]$
14:  $\mathbf{M}_{final},\mathbf{A}_{final},\mathbf{S}_{final}\leftarrow\text{Concat}\{(\mathbf{M}_{1},\mathbf{M}_{2}),(\mathbf{A}_{1},\mathbf{A}_{2}),(\mathbf{S}_{1},\mathbf{S}_{2})\}$
@@
16:  $\mathbf{B}_{raw}\leftarrow\text{BboxHead}(\mathbf{M}_{final},\mathbf{A}_{final})$
17:  $\mathbf{T}\leftarrow IoU_{N}+\mathbf{S}_{final}\times(IoU_{M}-IoU_{N})$
18:  $\mathbf{B}_{final}\leftarrow\text{DynamicNMS}(\mathbf{B}_{raw},\mathbf{S}_{final},\mathbf{T})$
@@
Secondly, attention is computed efficiently within each window, ensuring that object details are enhanced while minimizing redundant global interactions. To enhance cross-window communication, we introduce axis permuted attention, which enables not only region attentions but also long-range dependencies across high-confidence regions.
@@
Axis Permuted Encoder for Feature Enhancement. For each valid window, the corresponding feature map $F_{P}^{(i,j)}$ is added with corresponding relative and global positional encoding. Then the Axis Permuted Encoder (APE) refines feature representations by processing local windows through sequential self-attention operations while introducing a spatial permutation mechanism, as can be seen in Figure 4. Given the positional-encoded window feature $F_{W,k}^{(i,j)}$ where $k$ is the intermediate state index, the first self-attention pass is computed as follows:

### v2

@@ Abstract.
Tiny object detection plays a vital role in drone surveillance, remote sensing, and autonomous systems, enabling the identification of small targets across vast landscapes. However, existing methods suffer from inefficient feature leverage and high computational costs due to redundant feature processing and rigid query allocation. To address these challenges, we propose Dome-DETR, a novel framework with Density-Oriented Feature-Query Manipulation for Efficient Tiny Object Detection. To reduce feature redundancies, we introduce a lightweight Density-Focal Extractor (DeFE) to produce clustered compact foreground masks. Leveraging these masks, we incorporate Masked Window Attention Sparsification (MWAS) to focus computational resources on the most informative regions via sparse attention. Besides, we propose Progressive Adaptive Query Initialization (PAQI), which adaptively modulates query density across spatial areas for better query allocation. Extensive experiments demonstrate that Dome-DETR achieves state-of-the-art performance (+3.3 AP on AI-TOD-V2 and +2.5 AP on VisDrone) while maintaining low computational complexity and a compact model size. Code is available at https://github.com/RicePasteM/Dome-DETR.
@@
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3754861ccs: Computing methodologies Object detection
Figure 1. Comparisons with other detectors on the AI-TOD-V2 test set. More statistics is on Table 1.
@@ 1. Introduction
Object detection is a core task in computer vision, supporting applications like autonomous driving and robotic navigation. Recent progress has been largely driven by CNN-based methods (Dai et al., 2016; Redmon et al., 2016; He et al., 2017; Lin et al., 2017; Liu et al., 2016; Girshick, 2015; Ren et al., 2016; Wang et al., 2023; Tang et al., 2024; Peng et al., 2024b). More recently, Detection Transformers (DETR) introduced Transformer-based end-to-end detection (Carion et al., 2020). Subsequent variants (Zhu et al., 2020; Zhang et al., 2022; Peng et al., 2024a; Chen et al., 2024; Liu et al., 2022; Meng et al., 2021; Huang et al., 2024b; Zhao et al., 2024) have further enhanced performance and speed. However, these advancements primarily target generic object detection, leaving critical challenges unresolved for detecting tiny objects—particularly in aerial imagery from drones or satellites.
@@
Tiny object detection, which focuses on localizing and classifying objects occupying only a few pixels, is critical for tasks like remote sensing, drone monitoring, and autonomous navigation. Its challenge stems from the fragile and sparse feature representation of tiny objects. These objects heavily rely on low-level spatial details, which are often lost in deeper layers of feature hierarchies. Maintaining high-resolution feature maps help preserve such information but comes at the cost of increased computational complexity and memory consumption. Deformable DETR (Zhu et al., 2020) alleviates some of this burden through sparse deformable attention, yet still suffers from high inference latency due to its broad multi-scale attention. RT-DETR (Zhao et al., 2024) improves inference speed by decoupling intra-scale and cross-scale interactions but relies heavily on deep, low-resolution features, resulting in degraded performance on small-scale targets. These issues highlight the pressing need for detection frameworks that can balance fine-grained feature preservation with computational efficiency.
@@
Beyond feature representation, query allocation poses another challenge for tiny object detection, especially in aerial images with dense and complex object distributions (Wang et al., 2021b; Zhu et al., 2021). Existing DETR-like methods (Carion et al., 2020; Zhu et al., 2020; Liu et al., 2022; Li et al., 2022; Zhang et al., 2022; Zhao et al., 2024) use a fixed number of queries ($K=100$ in DETR, $K=300$ in Deformable-DETR), which simplifies implementation but reduces recall in dense scenes and wastes resources in sparse ones. Some aerial images (Wang et al., 2021b) contain over 1,500 tiny objects, far exceeding typical query capacities. DDQ-DETR (Zhang et al., 2023) increases query density (using $K=900$) and applies class-agnostic NMS with manually set IoU thresholds to filter redundant predictions. However, its rigid query count and fixed NMS thresholds introduce low recall in dense scenes and remain insensitive to variations in instance density. Meanwhile, DQ-DETR (Huang et al., 2024a) introduces dynamic query adjustment through a categorical counting module that estimates a classified number for query allocation. Although promising, its counting head relies on manually tuned classification hyperparameters across different datasets, and module design leads to substantial computational overhead. These challenges underscore the need for an adaptive query mechanism that dynamically aligns query density with instance distributions while eliminating manual tuning and preserving efficiency. To this end, tiny object detection remains challenging due to the fragile nature of small object features and the inefficiency of fixed-query mechanisms in handling diverse instance densities.
@@
Figure 2. (A) Comparison of foreground proportions across different datasets; and (B) Visualization of foreground regions (highlighted in yellow) in samples from these datasets.
In this paper, we introduce Dome-DETR, a novel and efficient end-to-end object detection framework which improves small and tiny object detection. First, in satellite, drone, and natural imagery, the foreground typically occupies only a small portion of the frame (see Figure 2), leading to excessive computation on less informative background regions. This highlights the potential for acceleration by emphasizing foreground areas. Moreover, shallow CNN features inherently capture rich spatial density cues. To leverage this, we design a lightweight (A) Density-Focal Extractor (DeFE) that produces density response heatmaps closely aligned with ground-truth annotations. These heatmaps enhance multi-scale encoder features and guide progressive query initialization in the decoder. Second, while deep features provide scene-level context, instance-level details in shallow features are essential for detecting small objects—but come with high attention costs. To mitigate this, we use DeFE-generated maps to suppress irrelevant regions via (B) Masked Window Attention Sparsification (MWAS), focusing attention on meaningful windows. Finally, fixed-query mechanisms in DETR struggle with large variations in object counts. We address this with (C) Progressive Adaptive Query Initialization (PAQI), which adaptively allocates queries by decoding density maps into dynamic suppression thresholds, removing the need for hand-tuned hyperparameters while improving recall in dense scenes.
@@
Experimental results on the AI-TOD-V2 (Wang et al., 2021b) dataset show that Dome-DETR achieves state-of-the-art performance in tiny object detection, outperforming existing models in both accuracy and efficiency. Specifically, Dome-DETR-M and Dome-DETR-L reach 34.0% (+3.2 AP) and 34.6% (+3.3 AP) on the AI-TOD-V2 test set, with only 252.6 and 358.7 GFLOPs, respectively. On the VisDrone (Zhu et al., 2021) val set, Dome-DETR-L also achieves 39.0% (+2.5 AP). By effectively tackling core challenges in feature representation and query allocation, Dome-DETR marks a substantial step forward for DETR-based tiny object detection. In summary, our main contributions are:
@@
  • We achieve state-of-the-art performance on the AI-TOD-V2 and VisDrone-DET-2019 datasets while maintaining a low computational cost, outperforming all existing models.
@@ 3.1. Overview
As shown in Figure 3, our study proposes Dome-DETR, which is built upon D-FINE (Peng et al., 2024a). We enhance the model with three components, i.e. (A) Density-Focal Extractor (DeFE) for density prediction, (B) Masked Window Attention Sparsification (MWAS) for efficient shallow feature enhancement, and (C) Progressive Adaptive Query Initialization (PAQI) for dynamic query manipulation.
@@ 3.2. Density-Focal Extractor
In DETRs, backbone generates spatially-channel encoded feature maps via multi-scale extraction, learned end-to-end to transform basic visual cues into high-level semantics. As a result, foreground object regions exhibit distinctive activations, while backgrounds maintain low responses. Prior studies show that shallow features retain vital spatial cues for small objects (Ning et al., 2023; Tu et al., 2022), but they are challenging to exploit due to: (1) High-resolution maps causing prohibitive computation, and (2) Redundant background areas dominating attention, weakening instance-level signals.
To tackle this, we propose the Density-Focal Extractor (DeFE), a lightweight module that explicitly learns instance density distributions to guide efficient feature enhancement and query allocation. As shown in Figure 3(A), DeFE processes the shallowest backbone feature map $F_{S}\in\mathbbm{R}^{H\times W\times C}$ through a cascaded network using depthwise separable convolutions with dilation rates (1, 2, 3) for multi-scale context capture while staying efficient. A lightweight attention block enhances salient areas via channel-wise recalibration. The output passes through a density head with a 3×3 convolution and bilinear upsampling, producing a normalized density heatmap $D_{pred}\in\mathbbm{R}^{H\times W\times 1}$. Thus, DeFE can be formulated as:
@@
where $f_{\mathrm{DSConv}}(\cdot)$ applies cascaded depthwise separable convolutions with varying dilation rates to extract spatially rich features. $F_{G}$ is a compact global representation obtained via Global Average Pooling (GAP). $W_{P}$ denotes a convolution layer projecting features into a single-channel density map, followed by a sigmoid function and bilinear upsampling to restore the original resolution and distribution.
The ground-truth density map $D_{gt}$ is generated by convolving Gaussian kernels centered at each object’s coordinates, with kernel sizes proportional to their bounding boxes. This encodes both instance locations and relative scales into a continuous supervision signal. To train DeFE, we propose Density Recall Focal Loss (DRFL), which emphasizes accurate density estimation in critical regions:
(4) $\hskip-5.0pt\mathcal{L}_{DRFL}=\sum_{i,j}^{H,W}\left[\alpha_{i,j}(d_{i,j}^{pred}-d_{i,j}^{gt})^{2}+\beta\cdot\mathbbm{1}(d_{i,j}^{pred}<d_{i,j}^{gt})\cdot d_{i,j}^{gt}\right],$
where $\alpha_{i,j}=\sqrt{d_{i,j}^{gt}}$ adaptively weights positions based on ground-truth density, and $\beta$ penalizes underestimation in dense regions. This ensures balanced learning across sparse and crowded areas while reducing FN in heavy occlusion. Despite high estimation accuracy, the module adds only 0.8M parameters thanks to optimized depthwise convolutions and attention.
3.3. Masked Window Attention Sparsification
The high resolution of shallow features is vital for capturing fine details of tiny objects but also incurs significant computational cost. Applying global attention to such dense maps results in high memory usage and latency. To alleviate this, we propose Masked Window Attention Sparsification (MWAS), which selectively focuses computation on key foreground regions while discarding redundant background. As shown in Figure 3(B), MWAS comprises two main stages: Foreground Token Pruning and Axis Permuted Encoder (APE).
Using the density map predicted by the DeFE, we first generate a binary mask that retains high-density regions and prunes low-density background. This significantly reduces the tokens involved in attention while preserving essential object details.
@@ 3.3. Masked Window Attention Sparsification
Density-Based Mask Generation. Given the shallow feature map $F_{S}$ from the backbone and the density heatmap $D_{\mathrm{pred}}\in\mathbbm{R}^{H\times W\times 1}$ from DeFE, where each element $d_{i,j}^{\mathrm{pred}}$ denotes the estimated object density at position $(i,j)$, we generate a binary mask $M_{b}\in{0,1}^{H\times W}$ by applying an adaptive threshold $T_{b}$ to $D_{\mathrm{pred}}$:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Secondly, attention is efficiently computed within each window, enhancing object details while minimizing redundant global interactions. To improve cross-window communication, we introduce axis permuted attention, enabling both regional focus and long-range dependencies across high-confidence areas.
@@
Axis Permuted Encoder for Feature Enhancement. For each valid window, the corresponding feature map $F_{P}^{(i,j)}$ is augmented with relative and global positional encodings. The Axis Permuted Encoder (APE) then refines features by processing these windows with sequential self-attention, incorporating a spatial permutation mechanism, as shown in Figure 4. Given the position-encoded window feature $F_{W,k}^{(i,j)}$, where $k$ is the intermediate state index, the first self-attention is computed as: